Question #2: A preparation for cough contains 1.5 grams of an expectorant per 100 mL. How many grain of the expectorant are there in a teaspoonful?
Answer by nerdybill(6948)

The key to this problem is to determine the volume of a single teaspoon (in metric units).
.
Reference:
http://www.joyofbaking.com/USMetricVolumeEquiv.html
.
1 teaspoon = 4.9 ml
.
Your proportion is then:
Let x = grams of expectorant in a teaspoon
1.5/100 = x/4.9
x = (1.5*4.9)/100
x = 7.35/100
x = .735 grams (or 735 milligrams)